OpenStudy (anonymous):

solve using the quadratic formula: x^2-6x+13=0

OpenStudy (anonymous):

i know it goes into\[\frac{ 6\sqrt{-6^2+4*1*13} }{ 2*1 }\]

OpenStudy (anonymous):

but I cant get the answer

OpenStudy (anonymous):

\[ \Delta=b^2-4 a c= 36 -52=-16 <0 \implies 2 \text {complex roots} \\ x=\frac { 6\pm i\sqrt{16}}{2}=3\pm 2i \\ \]

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

YW